OpenStudy (shealah):

What are two consecutive integers, such that five times the smaller minus four times the larger is 24?

OpenStudy (asnaseer):

do you know what "consecutive" means?

OpenStudy (shealah):

i know that it is x and x+1 but i just need to know how to set it up

OpenStudy (asnaseer):

good, so first number can be represented as x and next one as x+1 as you indicated. now the questions says that 5 times the smaller number minus 4 times the larger is equal to 24 so, how would you represent 5 times the smaller number?

OpenStudy (shealah):

5x

OpenStudy (asnaseer):

good, now how would you represent 4 times the larger number?

OpenStudy (shealah):

4(x+1)

OpenStudy (asnaseer):

great! now how would you use these two expressions to represent "5 times the smaller number minus 4 times the larger is equal to 24"?

OpenStudy (shealah):

5x - 4(x+1) = 24

OpenStudy (shealah):

would you do -4( x+1) or 4(x+1) then subtract?

OpenStudy (asnaseer):

your first attempt is correct

OpenStudy (asnaseer):

so now you just need to solve for x in: 5x - 4(x+1) = 24

OpenStudy (asnaseer):

if you type up your steps then I can check if you got them right

OpenStudy (shealah):

i got x to equal 20?

OpenStudy (asnaseer):

not quite, please show your steps so I can spot where you may have made a mistake

OpenStudy (shealah):

I think i got it this time... 5x - 4(x+1) = 24 5x-4x-4 = 24 x-4 = 24 x = 28

OpenStudy (asnaseer):

perfect! well done! :)

OpenStudy (shealah):

:)